Given x≡19(mod5)
Identify which one of the following options gives integers that are valid solutions for x ?


-19 and 7

-11 and 54

7 and 26

11 and 26


Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52754)

.
As formulated and presented, this post MAKES NO SENSE.


Answer by math_tutor2020(3816)


Answer: Choice B) -11 and 54

Reason

If a = b (mod n) then a-b is a multiple of n.
It means that a-b = nk for some integer k.

-11 = 19 (mod 5) because -11-19 = -30 is a multiple of 5.
54 = 19 (mod 5) because 54-19 = 35 is a multiple of 5.